Factor Weight East #1 East #2 West
Initial cost 10 130 130 100
Traffic 10 40 60 55
Maintenance 5 20 15 17
Dock space 6 18 17 24
Neighborhood 3 9 7 14


a.
Using the above factor ratings, calculate the composite score for each location.

Location Composite Score
East #1
East #2
West


b.
Determine which location has the highest composite score:

  • East #1

  • East #2

  • West
Answer #1

Answer:

A)

Composite Score for East 1 = (10*130+10*40+5*20+6*18+3*9)/(10+10+5+6+3)

Composite Score for East 1 =59.911

Composite Score for East 2 = (10*130+10*60+5*15+6*17+3*7)/(10+10+5+6+3)

Composite Score for East 2 =61.70

Composite Score for West = (10*100+10*55+5*17+6*24+3*14)/(10+10+5+6+3)

Composite Score for West =53.56

B)

From Above calculation we find that East#2 has the highest composite score.